OpenStudy (anonymous):

log4 1 - does it equal 4?

OpenStudy (anonymous):

\[\log _{4}(1)\] means 4 to what power equals one. The zero power of 4 equals 1 \[\log _{4}(1)_{ } = \log _{4}(4)^{0}_{ } \]

OpenStudy (anonymous):

log 4 1 ? what? log base 4 1 ?

OpenStudy (l):

It is equal to zero. *